IF sin 8 θ + cos 8 θ – 1 = 0, then what is the value of cos 2 θ sin 2 θ

 
IF sin8 θ + cos8 θ – 1 = 0, then what is the value of cos2 θ sin2 θ
 

Grade:11

1 Answers

Arun
5 years ago
Dear Omkar
 
sin8 @ + cos 8@ = 1
 
1/sqrt(2) * sin8@ + cos 8@ * 1/sqrt(2) = 1/sqrt(2)
 
Now 
sin (8@ + pi/4) = sin (pi/4)
 
8@ = 0 (only taking one solution )
@ = 0°
 
Hence
 
Sin2 @ * cos 2@ = sin 0°  * cos 0° = 0* 1 = 0